OpenStudy (unklerhaukus):

\[y^{\prime\prime}+4y^\prime=8x+6\sin2x\]

OpenStudy (unklerhaukus):

\[y_c^{\prime\prime}+4y_c^\prime=0\]\[\Delta =(4)^2-4(1)=12\]\[m^2+4m=0\]\[m=\frac{-4\pm2\sqrt3}{2}\]\[m_{1,2}=-2\pm\sqrt3\]\[y_c=Ae^{-2+\sqrt3x}+Be^{-2-\sqrt3x}\]

OpenStudy (unklerhaukus):

\[y_{p_1}^{\prime\prime}+4y_{p_1}^\prime=8x\]\[y_{p_1}=Cx+D;\qquad y^\prime_{p_1}=C:\qquad y^\prime_{p_1}=0\]\[0+4C=8x;\qquad C=2\]\[y_{p_1}=2x\]

OpenStudy (unklerhaukus):

\[y_{p_2}^{\prime\prime}+4y_{p_2}^\prime=6\sin2x\] \[y_{p_2}=E\sin (Fx)+G\cos(Hx)\]\[y_{p_2}^\prime=EF\cos(Fx) -GH\sin(Hx)\]\[ y_{p_2}^{\prime\prime}=-EF^2\sin(Fx)-GH^2\cos(Hx)\]\[\left(-EF^2\sin(Fx)-GH^2\cos(Hx)\right)+4\left(E\sin (Fx)+G\cos(Hx)\right)=6\sin2x\]\[(4-F^2)E\sin(Fx)+(4-H^2)G\cos(Hx)=6\sin2x\]

OpenStudy (unklerhaukus):

now what?

OpenStudy (unklerhaukus):

should i have \[F=H\]?

OpenStudy (turingtest):

you messed up at the beginning dude

OpenStudy (turingtest):

y''+4y'=0 so r^2+4r=0 the the characteristic polly for the complimentary

OpenStudy (unklerhaukus):

i see i should have\[\Delta=4^2\]

OpenStudy (turingtest):

I don't know what that delta notation is I see though that you do have m^2+4m=0 but you found the solution to be some crazy numbers

OpenStudy (unklerhaukus):

i was using delta for the i discriminant

OpenStudy (turingtest):

\[m^2+4m=0\implies m(m+4)=0; m=\{0,-4\}\]

OpenStudy (unklerhaukus):

that is right

OpenStudy (turingtest):

but you wrote something else above with radicals and stuff... anyway, so the complimentary is\[y_c=c_1+c_2e^{-4x}\]

OpenStudy (turingtest):

so for the guess for the particular wee need to add the guesses for the polynomial part and the trig part

OpenStudy (turingtest):

the guess for the polynomial part is\[Ax+B\]and the guess for the trig part is\[C\sin(2x)+D\cos(2x)\](you have to keep the 2x, it never changes with derivatives or integrals, right?)

OpenStudy (unklerhaukus):

ah good point

OpenStudy (turingtest):

so our guess for the particular will be\[Y_p=Ax+B+C\sin(2x)+D\cos(2x)\]

OpenStudy (turingtest):

oh wait, but I am missing something we only have y'' and y' so I think we need to add an extra x...

OpenStudy (unklerhaukus):

\[\Delta =(4)^2=16\]\[m^2+4m=0\]\[m(m-4)=0\]\[m_{1,2}=0,4\]\[y_c=A+Be^{4x}\] how did you get negative in the index

OpenStudy (turingtest):

how did you get from\[m^2+4m=0\]to\[m(m-4)=0\]o-0 ?

OpenStudy (unklerhaukus):

ah yes

OpenStudy (foolaroundmath):

I would suggest, instead of guessing the particular solution, try the annihilator method as given in these notes. They come in pretty handy to solve this sort of problems http://home.iitk.ac.in/~sghorai/TEACHING/MTH203/ode9.pdf

OpenStudy (turingtest):

so yeah, I missed that because we need to keep the solutions linearly independent, and we already have a constant as an answer in the particular, we need to multiply the polynomial part of our guess by x so it should be\[Y_p=Ax^2+Bx+C\sin(2x)+D\cos(2x)\]

OpenStudy (unklerhaukus):

\[y_{p_2}^{\prime\prime}+4y_{p_2}^\prime=6\sin2x\]\[y_{p_2}=E\sin (2x)+G\cos(2x)\]\[y_{p_2}^\prime=2E\cos(2x) -2G\sin(2x)\]\[ y_{p_2}^{\prime\prime}=-4E\sin(2x)-4G\cos(2x)\]

OpenStudy (turingtest):

ok...

OpenStudy (unklerhaukus):

is annihilator method the same as using operator D?

OpenStudy (foolaroundmath):

yes

OpenStudy (turingtest):

I am self-taught in DE's, so I don't know that terminology :(

OpenStudy (unklerhaukus):

that is amazing turing test

OpenStudy (foolaroundmath):

I concur with Unkle :O. I certainly wouldnt have been able to self teach DE's. I'm still not familiar with PDE's

OpenStudy (turingtest):

thanks, I'm reading this annihilator thing now I've seen these D's but never understood what they mean

OpenStudy (unklerhaukus):

\[\left(-4E\sin(2x)-4G\cos(2x)\right)+4\left(E\sin (2x)+G\cos(2x)\right)=6\sin2x\] \[-3G\cos(Hx)=6\sin2x\]

OpenStudy (unklerhaukus):

how did that H get in there ....H=2

OpenStudy (unklerhaukus):

\[G=-2\tan (2x)\] have i made more mistakes yet?

OpenStudy (fwizbang):

In this particular problem, it would seem to make a lot a sense to simply integrate both sides before doing anything(since the RHS is a derivative.). Then you can apply your favorite method to the resulting first order eq.....

OpenStudy (unklerhaukus):

why diden't i see that earlier @fwizbang .

OpenStudy (turingtest):

this isn't going to work look Unk, you did mess up, everything cancels, even the cos

OpenStudy (unklerhaukus):

hmm

OpenStudy (unklerhaukus):

-4+4≠3?

OpenStudy (turingtest):

oh wait, you put y for y'

OpenStudy (turingtest):

shoud be\[\left(-4E\sin(2x)-4G\cos(2x)\right)+4\left(2E\cos (2x)+2G\sin(2x)\right)=6\sin2x\]

OpenStudy (turingtest):

-sin I mean..

OpenStudy (unklerhaukus):

i did too, how embarrassing, failure to substitute correctly on my part

OpenStudy (turingtest):

\[\left(-4E\sin(2x)-4G\cos(2x)\right)+4\left(2E\cos (2x)-2G\sin(2x)\right)=6\sin2x\]

OpenStudy (unklerhaukus):

\[\left(-4E\sin(2x)-4G\cos(2x)\right)+4\left(2E\cos(2x) -2G\sin(2x)\right)=6\sin(2x)\]

OpenStudy (unklerhaukus):

good they match ,

OpenStudy (foolaroundmath):

Short and sweet using Annihilator's .. Basically \(D \equiv d/dx\) For the particular solution, \(y_{p}\), it must satisfy \((D^{2}+4D)y_{p} = 8x+6sin2x\) Now to annihilate \(8x\), we operate \(D^{2}\) and to annihilate \(\sin{2x}\) we operate \(D^{2}+4\), so operating \(D^{2}(D^{2}+4) \) with the DE, we get \( D^{3}(D+4)(D^{2}+4)y_{p} = 0\) this gives \(y_{p} = c_{1} + c_{2}x + c_{3}x^{2} + c_{4}e^{-4x}+c_{5}\sin{2x}+c_{6}\cos{2x}\) Since, \(c_{1} + c_{4}e^{-4x} \) form part of the homogenous solution, therefore the particular solution \(y_{p} = c_{2}x + c_{3}x^{2}+c_{5}\sin{2x}+c_{6}\cos{2x} \)

OpenStudy (unklerhaukus):

ill try annihilation with operator D after i know i can finish with equating co-efficient

OpenStudy (turingtest):

I don't get the D^2+4 to annihilate 6sin(2x) could someone explain that?

OpenStudy (turingtest):

oh wait I think I get it...

OpenStudy (turingtest):

\[\frac{d^2}{dx^2}(6\sin2x)+4(6\sin(2x)=0\]is that the idea?

OpenStudy (turingtest):

cool

OpenStudy (foolaroundmath):

The annihilator in essence is related to the solution of the characteristic equations. Note that sin(2x) will occur as a solution only if the roots of the characteristic equation are \(\pm 2i\). i.e. the characteristic equation is \(m^{2}+4\), consequently meaning that if we operate \{D^{2}+4\) on \(\sin{2x}\) we get 0.

OpenStudy (foolaroundmath):

yes @TuringTest

OpenStudy (turingtest):

gotchya, thanks :) I get it now

OpenStudy (unklerhaukus):

\[\left(-4E\sin(2x)-4G\cos(2x)\right)+4\left(2E\cos(2x) -2G\sin(2x)\right)=6\sin(2x)\] \[-4(E+G)\sin(2x)+4(2E-G)\cos(2x)=6\sin(2x)\] have i got this step right?

OpenStudy (unklerhaukus):

ah no i see an other mistake already

OpenStudy (unklerhaukus):

the first G in the second line should have a 2 in front of it,

OpenStudy (unklerhaukus):

\[-4(E+2G)\sin(2x)+4(2E-G)\cos(2x)=6\sin(2x)\] what should my next move be

OpenStudy (turingtest):

oh, set the coefficients equal\[-4(E+2G)=6\]\[4(2E-G)=0\]and solve the system

OpenStudy (unklerhaukus):

so from the your second equation G=2E

OpenStudy (turingtest):

I guess, yeah do you see where I got these equations from?

OpenStudy (unklerhaukus):

yeah because these are the coefficients of the trig functions , and must be equal

OpenStudy (turingtest):

yeah

OpenStudy (unklerhaukus):

\[G=-\frac{3}{10}\]?

OpenStudy (turingtest):

erm... either G or E is -3/10 I'm getting a little confused :S

OpenStudy (turingtest):

yeah I get E=-3/10

OpenStudy (unklerhaukus):

ok,

OpenStudy (unklerhaukus):

you were right i did get E/G mixed up

OpenStudy (unklerhaukus):

\[E=-\frac{3}{10};\qquad G=-\frac 35\]

OpenStudy (turingtest):

yep, so now we move on to finding the values of D and E (or whatever you called them) for the other particular

OpenStudy (turingtest):

oh you may want to write out this particular in full just so you don't forget which coefficient goes where

OpenStudy (unklerhaukus):

\[y_{p_2}=-\frac 3{10}\sin (2x)-\frac5{10}\cos(2x)\]

OpenStudy (turingtest):

yep, so now we do \(y_{p1}\)

OpenStudy (unklerhaukus):

i think i have \( y_{p_1} \) in my second post for this question

OpenStudy (unklerhaukus):

oh but it is wrong

OpenStudy (turingtest):

it's wrong because the solution you guessed is not linearly independent from the complimentary, as I mentioned earlier

OpenStudy (unklerhaukus):

hmm i have made so many mistakes

OpenStudy (turingtest):

we already know the complimentary is\[y_c=Ae^{-4x}+B\]so if\[y_p=Cx+D\]then we have two constants, B and D, hence the solutions are not linearly independent to fix this problem we need to multiply our guess for \(y_p\) by x

OpenStudy (turingtest):

\[y_{p1}=Cx^2+Dx\]now our complimentary and particular solutions are linearly independent, as they should be

OpenStudy (turingtest):

"if you don't make mistakes it means you didn't try hard enough" is a paraphrase of a favorite quote of mine

OpenStudy (unklerhaukus):

so we are multiplying by x because the original equation is \(y\)-absent?

OpenStudy (unklerhaukus):

\[y_{p_1}=Cx^2+Dx;\qquad y^\prime_{p_1}=2Cx+D:\qquad y^{\prime\prime}_{p_1}=2C\] \[2C+4Cx^2+Dx=8x;\qquad D=\frac 12\]

OpenStudy (turingtest):

we need the solutions to be linearly independent in other words we have to keep the wronskian from being zero if we allow one of our solutions to be linearly independent on the other (the constant solution in the complimentary, A is linearly dependent on part of our guess for the particular, Cx+D) so this would not be a fundamental set of solutions multiplying by x fixes that problem

OpenStudy (turingtest):

you could say that it is "because we don't have a y" in this particular problem, but a better way to think of it is that our guess for the particular cannot have any linear dependence on the complimentary solution, so when we see "oh, I have a constant in my complimentary and a constant for my guess for the particular, I need to fix that" no part of the complimentary should be a constant multiple of the particular

OpenStudy (turingtest):

and D should be -1/2

OpenStudy (turingtest):

you forgot to multiply D by the 4

OpenStudy (unklerhaukus):

hahah

OpenStudy (turingtest):

actually, a lot of that is wrong; I think you are being lazy, distracted or tired... I've seen you do much harder stuff\[y''+4y'=2C+8Cx+4D=8x\]

OpenStudy (unklerhaukus):

yeah its 417am, i am not tired enough to be asleep

OpenStudy (turingtest):

that would explain it look over this when you have your wits about you http://tutorial.math.lamar.edu/Classes/DE/UndeterminedCoefficients.aspx

OpenStudy (unklerhaukus):

turns out im not very at substitution

OpenStudy (turingtest):

lol

OpenStudy (unklerhaukus):

\[y_{p_1}^{\prime\prime}+4y_{p_1}^\prime=8x\] \[y_{p_1}=Cx^2+Dx;\qquad y^\prime_{p_1}=2Cx+D:\qquad y^{\prime\prime}_{p_1}=2C\] \[2C+8Cx+4D=8x\]\[2(C+2D)=0;\qquad 8Cx=8x\]\[C=1;\qquad D=-\frac 12\]

OpenStudy (unklerhaukus):

so \[y_c=A+Be^{-4x}\]\[y_{p_1}=x^2-\frac x2\]\[y_{p_2}=-\frac{3}{10}\left(\sin 2x+2\cos2x\right)\] \[y=y_c+y_{p_1}+y_{p_2}\] \[y(x)=A+Be^{-4x}+x^2 -\frac x2-\frac{3}{10}\left(\sin 2x+2\cos2x\right)\]

OpenStudy (unklerhaukus):

all done , fwizbang get the prize for simplest solution